# Nonlinear Fore(Back)casting and Innovation Filtering for Causal-Noncausal VAR Models

### v3

@@
Abstract We introduce closed-form formulas of out-of-sample predictive densities for forecasting and backcasting of mixed causal-noncausal (Structural) Vector Autoregressive VAR models. These nonlinear and time irreversible non-Gaussian VAR processes are shown to satisfy the Markov property in both calendar and reverse time. A post-estimation inference method for assessing the forecast interval uncertainty due to the preliminary estimation step is introduced too. The nonlinear past-dependent innovations of a mixed causal-noncausal VAR model are defined and their filtering and identification methods are discussed. Our approach is illustrated by a simulation study, and an application to cryptocurrency prices.
Keywords: Predictive Density, Backcasting, Filtering, Nonlinear Innovations, Generalized Covariance (GCov) Estimator, Cryptocurrency Bubble
@@ 1 Introduction
There has been a growing interest in applications involving the stationary mixed causal-noncausal Vector Autoregressive (VAR) models and their theoretical properties [Gourieroux, Jasiak (2017),(2022), Davis, Song (2020), Lanne, Luoto (2016),(2021), Swensen (2022), Cubbada, Hecq, Voisin (2023)]. In applied research, the stationary mixed VAR models can replicate local trends, spikes and bubbles characterizing, e.g. the commodity prices and cryptocurrency rates, which are interpreted as nonstationary features in the traditional literature on causal, i.e. past dependent processes. In terms of theoretical properties, the main differences between the mixed VAR’s and the traditional causal VAR’s are in the assumptions concerning the eigenvalues of the autoregressive matrix coefficients and the errors of the model. More specifically, in the mixed model the roots of the autoregressive characteristic equation can be in modulus either greater or smaller than 1, as opposed to being only greater than 1 in the causal VARs. While the errors of a causal VAR are assumed to be normally distributed, the errors of a mixed VAR need to be non-Gaussian for the identification of causal and noncausal dynamics. Thus, the traditional Gaussian Maximum Likelihood (ML) and Least Squares estimators are flawed in applications to mixed VAR’s. Under parametric distributional assumptions, the ML estimators based on non-Gaussian likelihood functions can be used for the mixed VAR models, as shown in Davis, Song (2020). When the error distribution is left unspecified, the Generalized Covariance (GCov) estimator is available, providing consistent, asymptotically normally distributed and semi-parametrically efficient parameters estimates in one step [Gourieroux, Jasiak (2023)].
There is an important difference between the errors of causal and mixed VAR models, other than non-Gaussianity. Unlike the errors of causal VARs, those of mixed VAR’s are not uncorrelated / independent from the past values of the process. Hence, they cannot be interpreted as innovations. For this reason, innovation-based inference, such as impulse response analysis (IRF) and forecast error variance decomposition (FEVD) routinely conducted in the causal VAR models cannot be applied in the same way to the mixed VAR models  [^3] See, e.g. Lutkepohl (1990).. It has to account for the nonlinear dynamic features of the mixed VAR model. So far, the nonlinear causal innovations of a mixed VAR model have not been defined in the literature.
Another difficulty arises with the fore- and backcasting of the mixed VAR models. A mixed VAR model with non-Gaussian errors cannot be forecast from its conditional expectation only, unlike the traditional causal VAR model. For out-of-sample (oos) nonlinear forecasting of mixed causal-noncausal VAR processes, there exist in the literature a simulation-based method of forecasting given in Nyberg, Saikkonen (2014) and a Bayesian method in Lanne, Luoto (2016), which are valid for a constraint multiplicative mixed VAR representation. An operational causal predictive density in closed-form for the multivariate mixed VAR models has been deemed infeasible in Fries, Zakoian (2019) 444”The predictive density is generally not available under closed form”, Fries, Zakoian (2019).
Moreover, the backcasting, which is straightforward for the time reversible linear Gaussian causal VAR processes, is complicated in this context as well. The mixed VAR does not satisfy the assumption of Gaussianity ensuring the time reversibility, which underlies the commonly used time series and machine learning methods of backcasting [Twumasi, Twumasi (2022)]. A closed-form formula of nonlinear backcasting for mixed VAR models has not been introduced in the literature yet.
This paper addresses the two problems of i) the lack of closed-form predictive densities for forecasting and backcasting and ii) the lack of a definition of nonlinear causal innovations for the mixed VAR model and their use for nonlinear impulse response analysis.
We provide closed-form formulas of backward and forward predictive density for out of sample forecasting and backcasting of mixed causal-noncausal VAR processes. For fore(back)casting, we show that the mixed VAR is a nonlinear Markov process in both calendar and reverse time while being time-irreversible. The closed-form expression of predictive densities for fore- and backcasting the mixed causal-noncausal VAR processes are given at horizon 1. They can be used sequentially to predict at any horizon.
From the predictive densities, we infer the point forecasts and prediction intervals. A post-estimation inference method for assessing the forecast interval uncertainty due to the preliminary estimation step is introduced too. A confidence set of the predicted set is proposed as a post-estimation inference method to capture the effect of the preliminary estimation step on the prediction interval.
The nonlinear causal innovations are defined by extending the results in Koop, Pesaran, Potter (1996), Potter (2000), and Gourieroux, Jasiak (2005). We determine their relationship with the model errors and discuss their identification and in-sample filtering.
The forecast performance and the nonlinear causal innovations analysis are illustrated by simulations and empirical applications. We use the one-step semi-parametric GCov estimators introduced in Gourieroux, Jasiak (2023) for VAR models. This approach does not require any distributional assumptions on the errors of the model, except for their non-Gaussianity. The semi-parametric estimation method distinguishes our approach from the existing literature employing maximum likelihood-based methods [Andrews et al. (2006), Davis, Song (2020), Swensen (2022)], which are consistent provided that the distributional assumptions are satisfied.
@@
The paper is organized as follows. Section 2 describes the causal-noncausal VAR model and its constrained multiplicative form. Section 3 recalls the state-space representation of the mixed VAR model and derives the closed-form formulas of the multivariate predictive density for forecasting and backcasting. Section 4 defines the nonlinear causal innovations and discusses their identification and filtering. Section 5 reviews the semi-parametric estimation of the mixed VAR model and introduces a filtering algorithm and post-estimation inference on the random prediction set. A simulation study and an empirical application to a bivariate process of Bitcoin/USD and Ethereum/USD exchange rates are provided in Section 6. Section 7 concludes. The technical results are given in Appendices A.1-A.3. Appendix A.1 contains the proof of the forward and backward predictive density formula. Appendix A.2 explains the constraints induced by the multiplicative representation of a causal-noncausal VAR model. Online Appendix A.3 examines the (under)-identification of nonlinear innovations in a multivariate framework.
Additional information is provided in the online Appendices B, C and D, which include a discussion of the identification conditions for Independent Component Analysis (ICA), the closed-form expression of the kernel-based semi-nonparametric estimator of predictive density, and additional results for the empirical applications.
@@ 2 Mixed Causal-Noncausal Processes
This Section reviews the causal-noncausal VAR(p) model studied in Gourieroux, Jasiak (2017), Davis, Song (2020) 555based on a working paper from 2012., and Swensen (2022).
@@ 2.1 The Model
where $Y_{t}$ is a vector of size $m$, $\Phi_{j},\,j=1,...,p$ are matrices of autoregressive coefficients of dimension $m\times m$ and $(\varepsilon_{t})$ is a sequence of errors, which are serially independent, identically distributed (i.i.d.) random vectors of dimension $m$ with mean zero and finite variance-covariance matrix $\Sigma$. Errors $(\varepsilon_{t})$ are assumed to have a non-Gaussian distribution. Since $(\varepsilon_{t})$ is not assumed independent of past $Y$’s, this error process cannot be interpreted as an innovation process.
We assume that the roots of the characteristic equation of the autoregressive polynomial matrix $det(Id-\Phi_{1}\lambda-\cdots\Phi_{p}\lambda^{p})=0$ are of modulus either strictly greater, or strictly less than one, i.e. are either outside, or inside the unit circle. Then, there exists a unique strictly stationary solution to the VAR model (2.1).
The strictly stationary solution $(Y_{t})$ to model (2.1) can be written as a two-sided moving average in errors $\varepsilon_{t}$:
@@
This is a linear time series, according to the terminology of Rosenblatt (2012). The matrices of coefficients on the past and future terms of this MA representation are uniquely defined when $\varepsilon_{t}$ is non-Gaussian, which is an identifying assumption. Process $Y_{t}$ is said to be causal in $\varepsilon_{t}$, if $Y_{t}=\sum_{j=0}^{+\infty}C_{j}\varepsilon_{t-j}$, noncausal in $\varepsilon_{t}$, if $Y_{t}=\sum_{j=-\infty}^{-1}C_{j}\varepsilon_{t-j}=\sum_{j=1}^{+\infty}C_{-j}\varepsilon_{t+j}$, or mixed, otherwise.
In the presence of a noncausal component, the assumption of strict stationarity of ($Y_{t}$) implies the nonlinear causal dynamics of $Y_{t}$ with past-dependent conditional heteroscedasticity. Then, the process $(Y_{t})$ can be characterized by a possibly complicated conditional distribution of $Y_{t+h}$ given $\underline{Y_{t}}=(Y_{t},Y_{t-1},...)$ for $h=1,2,..$, that provides nonlinear out-of-sample (oos) predictions.
A process $(Y_{t})$ with nonlinear causal dynamics may display local trends, spikes and bubbles similar to those observed in the time series of commodity (oil) prices, exchange rates, or cryptocurrency prices [Gourieroux, Zakoian (2017), Gourieroux, Jasiak (2017), Gourieroux, Jasiak, Tong (2021)].
2.2 Multiplicative VAR
A constrained multiplicative Mixed Autoregressive (MAR) representation of the mixed VAR model was proposed by Lanne, Saikkonen (2013): $\Phi(L)\Psi(L^{-1})Y_{t}=\epsilon_{t}^{*},$ where the matrices of autoregressive polynomials $\Phi$ and $\Psi$ have both roots outside the unit circle and $\epsilon_{t}^{*}$ is an i.i.d. sequence of errors. Davis, Song (2020), Swensen (2022), point out that the multiplicative representation of a mixed causal-noncausal VAR model does not always exist and Cubbada, Hecq, Voisin (2022) note that it is not equivalent to $\Psi(L^{-1})\Phi(L)Y_{t}=\epsilon_{t}^{*}$, since $\Psi(L^{-1})$ and $\Phi(L)$ do not commute, in general. More precisely, the multiplicative representation exists for the univariate mixed processes [Breidt et al. (1991), Lanne, Saikkonen (2008)] and for the multivariate mixed VAR(1) process [Gourieroux, Jasiak (2017), Corollary 3 and Section 5]. Otherwise, Appendix A.2 shows that the multiplicative representation may not be compatible with the representation (2.1) of the VAR model. For example, the multiplicative VAR(2) model $(Id-\Phi L)(Id-\Psi L^{-1})Y_{t}=\varepsilon_{t}^{*}$ where the matrix polynomials have roots of modulus strictly greater than 1 and $(\varepsilon_{t}^{*})$ are i.i.d., can be written under the form $Y_{t}=\Phi_{1}Y_{t-1}+\Phi_{2}Y_{t-2}+\varepsilon_{t}$. However, the errors $\varepsilon_{t}$ are independent only if the matrix $\Psi$ is invertible.
The set of multiplicative VAR and mixed VAR models are non-nested, i.e. a mixed model does not always admit a multiplicative form and a multiplicative VAR model cannot always be written as a mixed VAR. We focus our attention on the mixed VAR model in this paper.
3 State-Space Representation and Predictive Density
Let us recall the state-space representation of the mixed VAR process.
@@ 3.1 State-Space Representation
We consider a mixed VAR(p) process of dimension $m$:
$Y_{t}=\Phi_{1}Y_{t-1}+\cdots+\Phi_{p}Y_{t-p}+\varepsilon_{t},$ (3.1)
where the serially i.i.d. error vectors $\varepsilon_{t}$ have a continuous joint distribution with probability density function $g$. These parameters are assumed given in this Section. We assume that the roots of:
$\det(Id-\Phi_{1}z-\cdots-\Phi_{p}z^{p})=0,$ (3.2)
are not on the unit circle. Then, there exists a unique stationary solution of model (3.1) with a two-sided moving average representation in $(\varepsilon_{t})$.
@@
$\left(\begin{array}[]{c}Y_{t}\\ \tilde{Y}_{t-1}\end{array}\right)\equiv\left(\begin{array}[]{c}Y_{t}\\ Y_{t-1}\\ \vdots\\ Y_{t-p+1}\end{array}\right)=\Psi\left(\begin{array}[]{c}Y_{t-1}\\ \tilde{Y}_{t-2}\end{array}\right)+\left(\begin{array}[]{c}\varepsilon_{t}\\ 0\end{array}\right),$ (3.3)
with
$\Psi=\left(\begin{array}[]{cccc}\Phi_{1}&\cdots&\cdots&\Phi_{p}\\ Id&0&\cdots&\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&Id&0\end{array}\right).$ (3.4)
The eigenvalues of the autoregressive matrix $\Psi$ are the reciprocals of the roots of the characteristic equation (3.2).
@@
Matrix $\Psi$ has a real Jordan representation: $\Psi=A\left(\begin{array}[]{cc}J_{1}&0\\ 0&J_{2}\end{array}\right)A^{-1},$ where $J_{1}$ (resp. $J_{2}$) are real $(n_{1}\times n_{1})$ (resp. $(n_{2}\times n_{2})$) matrices where $n_{2}=n-n_{1}$ with all eigenvalues of modulus strictly less than 1 [resp. strictly larger than 1] and $A$ is a ($n\times n$) invertible matrix [see Perko (2001), Gourieroux, Jasiak (2016), Section 5.2, Davis, Song (2020), Remark 2.1. for real Jordan representations]. Then, equation (3.3) can be rewritten after the change of basis $A^{-1}$ as:
@@
Let us introduce a block decomposition of $A^{-1}$:
$A^{-1}\equiv\left(\begin{array}[]{c}A^{1}\\ A^{2}\end{array}\right),$ (3.6)
where $A^{1}$ is of dimension $(n_{1}\times n)$ and the transformed variables:
@@
Then, we get two sets of state components of process ($Y_{t}$) such that:
@@
The state-space representation of process ($Y_{t}$) consists of:
state equations of state variables $Z_{t}$ that follow the causal and non-causal equations (3.8);
measurement equations obtained by solving for $Y_{t}$ from: $\left(\begin{array}[]{c}Y_{t}\\ \tilde{Y}_{t-1}\end{array}\right)=AZ_{t}.$
These measurement equations are deterministic and the filtrations generated by $(Y_{t})$ and $(Z_{t})$ are identical.
Since the Jordan representation is not unique, the state-space representation is not unique either. It depends on the choice of state variables, i.e. the latent factors $Z_{1}$ and $Z_{2}$, up to linear invertible transformations.
c) State Linear Innovations
The first set of state equations (3.8) defines the causal VAR(1) process ($Z_{1,t}$). Thus, the process $(Z_{1,t})$ has the causal MA()($\infty$)( ∞ ) representation:
@@
with $\eta_{1,t}$ as a function of $\varepsilon_{t},\varepsilon_{t-1}...$.
The second set of state equations (3.8) needs to be inverted to obtain a MA representation in matrices with eigenvalues of modulus strictly less than 1. We get:
$Z_{2,t}=J_{2}^{-1}Z_{2,t+1}-J_{2}^{-1}\,\eta_{2,t+1},=-\sum_{j=0}^{+\infty}[J_{2}^{-j-1}\,\eta_{2,t+j+1}].$ (3.10)
Therefore, $(Z_{2,t})$ is a noncausal process with a one-sided moving average representation in future values $\varepsilon_{t+1},\varepsilon_{t+2},...,$.
Let us now discuss the state-specific linear innovations $\eta_{1,t},\eta_{2,t}$. From equations (3.9), (3.10) and the stationarity conditions, it follows that:
@@
where $\underline{Z}_{1,t-1}=Z_{1,t-1},Z_{1,t-2},...$ and $\bar{Z}_{2,t+1}=Z_{2,t+1},Z_{2,t+2},...$. Therefore, $\eta_{1,t}$ (resp. $\eta_{2,t}$) can be interpreted as the causal linear innovations of $Z_{1,t}$ based on the information $\underline{Z}_{1,t-1}$ (resp. noncausal linear innovation of $Z_{2,t}$ based on the information $\bar{Z}_{2,t+1}$). It is important to note that the information set $\underline{Z}_{1,t-1}$ (resp. $\bar{Z}_{2,t+1}$) differs in general from the global causal information set $\underline{Y}_{t-1}=\underline{Z}_{t-1}$ (resp $\bar{Y}_{t+1}=\bar{Z}_{t+1}$). Later, in Section 4, we clarify this interpretation by introducing the notion of a nonlinear innovation that takes into account all available information. To do that we need first to derive the expressions of forward and backward predictive densities.
3.2 Out-of-Sample Predictive Density
The expression of the predictive density of $Y_{T+1}$ given $\underline{Y}_{T}=(Y_{T},Y_{T-1},...)$ is given below and derived in Appendix A.1. This expression is written for given $\Phi_{1},\Phi_{2},...$ and joint error density $g$.

### v4

@@
Abstract
We show that the mixed causal-noncausal Vector Autoregressive (VAR) processes satisfy the Markov property in both calendar and reverse time. Based on that property, we introduce closed-form formulas of forward and backward predictive densities for point and interval forecasting and backcasting out-of-sample. The backcasting formula is used for adjusting the forecast interval to obtain a desired coverage level when the tail quantiles are difficult to estimate. A confidence set for the prediction interval is introduced for assessing the uncertainty due to estimation. We also define new nonlinear past-dependent innovations of mixed causal-noncausal VAR models for impulse response function analysis. Our approach is illustrated by simulations and an application to oil prices and real GDP growth rates.
Keywords: Mixed Causal-Noncausal Process, Bubble, Predictive Density, Backcasting, Nonlinear Innovations, Generalized Covariance (GCov) Estimator, Oil Price.
@@ 1 Introduction
The causal-noncausal (mixed) Vector Autoregressive processes have nonlinear dynamics with locally explosive patterns such as local trends, spikes and bubbles [Gourieroux and Jasiak (2017), Davis and Song (2020)]. In applied research, these strictly stationary models can replicate the behavior of commodity prices, including oil prices [Lof and Nyberg (2017), Cubadda et al. (2023), Blasques et al. (2025)], cryptocurrency rates [Hencic and Gourieroux (2019), Hall and Jasiak (2024), Cavaliere et al. (2020)], financial indexes, including S&P500 and NASDAQ [Gourieroux, Zakoian (2017), Freis (2022)], climate risk on El Nino and La Nina [De Truchis, Fries and Thomas (2024)], and green stocks prices [Hecq et al (2024)]. Specifically, a mixed VAR(1) process ($Y_{t}$) of dimension $m$ is the strictly stationary solution of $Y_{t}=\Phi Y_{t-1}+\varepsilon_{t}$, where the errors $\varepsilon_{t}$ are assumed to be non-Gaussian, independent and identically distributed (i.i.d.), and the eigenvalues of of autoregressive matrix $\Phi$ are of modulus either smaller or greater than 1, and are referred to as the causal and non-causal eigenvalues, respectively. In the presence of non-causal eigenvalues, the errors $\varepsilon_{t}$ are correlated with the lagged values of $Y_{t}$ and cannot be interpreted as the (causal) innovations of the process. We show that the mixed VAR model has a causal, i.e. past-dependent representation, which is a multivariate non-linear autoregression $Y_{t}=a(Y_{t-1};v_{t})$, where $a$ is a non-linear function, and the errors $v_{t}$ are i.i.d. and independent of $Y_{t-1}$, satisfying the definition of causal innovations.
The initial papers on mixed VARs were focused on the identification of noncausal dynamics and parameter estimation. Davis and Song (2020) introduced the maximum likelihood estimator under a parametric assumption on the distribution of error $\epsilon_{t}$. Gourieroux and Jasiak (2017), (2023) proposed a semi-parametric one-step Generalized Covariance (GCov) estimator which is consistent and semi-parametrically efficient. This paper is focused on the Markov property of mixed VAR processes, which leads to nonlinear predictive density formulas and nonlinear impulse response functions (IRF) for these processes. We derive the closed-form expressions of forward and backward predictive densities, and build prediction intervals (sets) that account for estimation risk and are adjustable to the desired conditional coverage. We also discuss the identification of the aforementioned causal innovations and define the nonlinear IRF for tracing out the effects of shocks to the locally explosive component of the process. The aim of this paper is to provide a toolbox of inference methods for mixed VAR models, equivalent to those available for the traditional causal VARs, to facilitate the use of these models in applied research in finance and macroeconomics.
The closed-form predictive density for forecasting the mixed VAR processes is new and derived in a semi-parametric framework. This closed-form representation was deemed infeasible until recently333”The predictive density is generally not available under closed form”, Fries and Zakoian (2019). It extends to a general framework earlier results on a mixed VAR model with a multiplicative form of autoregressive matrix, also called the multivariate Mixed Autoregressive (MAR) model [see Nyberg and Saikkonen (2014), Gourieroux and Jasiak (2016), Lanne and Luoto (2016)]. The two types of models are not equivalent, except for pure causal and noncausal processes, because the multiplicative form assumption is strongly constraining and may not exist for all mixed VAR processes. In particular, it does not exist for a bivariate, mixed VAR(1) model with at least one causal and one non-causal eigenvalues [see Appendix A.2]. The availability of a closed-form predictive density for the general VAR model considered in this paper eliminates the need for assuming a restrictive multiplicative form of autoregressive matrix and using computationally intense forecasting methods based on numerical approximations.
From the quantiles of the predictive densities, we infer the point forecasts and prediction intervals. In addition, we derive a new backcasting method for noncausal processes. The backcasting algorithms are readily available for time-reversible linear Gaussian time series, but need to be developed specifically for the time-irreversible non-Gaussian mixed VAR processes for the treatment of missing data, or implementation of backpropagation algorithms in applied research [Twumasi and Twumasi (2022)]. In our paper, we use the backcasting algorithm to build a new bootstrapped prediction interval adjustable to a desired coverage level, and we show how to account for the estimation risk by introducing the confidence set for prediction intervals. This approach is applicable to any nonlinear dynamic model where the forecast error is not a sum of the unknown future innovation and estimation error, and the process is back-castable. The adjustment of the empirical coverage of a prediction interval to the desired one is beneficial when the multivariate predictive density is estimated from data with a limited number of tail observations, or evaluated over insufficiently support points because of a computational burden.
Another contribution of this paper is in defining the nonlinear causal innovations for mixed VAR models and examining their identification issues. We describe the nonlinear IRF analysis of mixed VARs, with a shock applied to a locally explosive component of the process, and performed on- and off-bubble. The nonlinear causal innovation for the mixed VAR model is based on Gourieroux and Jasiak (2005), which has recently received attention in the context of nonlinear impulse response functions in macroeconomics [see, e.g. Gonzalves et al.(2021), Gourieroux and Lee (2025)].
@@
The paper is organized as follows. Section 2 describes the causal-noncausal VAR model and its state-space representation. Sections 3, 4 and 5 contain the new results. Section 3 proves the Markov property and derives the closed-form formulas of multivariate predictive density for forecasting and backcasting. Section 4 introduces the inference on the random set of prediction intervals. Section 5 defines the nonlinear causal innovations and discusses their identification and filtering. A simulation study and an empirical application to a bivariate process of oil prices and real US GDP rates are presented in Section 6. Section 7 concludes. The technical results are given in Appendices A.1-A.2. Appendix A.1 contains the proof of the forward and backward predictive density formula. Appendix A.2 explains the constraints induced by the multiplicative representation of a causal-noncausal VAR model. Online Appendices B, C, D and E present the closed-form expression of a kernel-based semi-nonparametric estimator of predictive density, give additional results on simulations and the empirical application of the mixed VAR model to oil prices and real GDP rates.
@@ 2 Mixed Causal-Noncausal Processes
This Section reviews the state space representation of the causal-noncausal (mixed) VAR(p) model studied in Gourieroux and Jasiak (2016),(2017), and Davis and Song (2020).
@@ 2.1 The Model
where $Y_{t}$ is a vector of size $m$, $\Phi_{j},\,j=1,...,p$, are matrices of autoregressive coefficients of dimension $m\times m$ and $(\varepsilon_{t})$ is a sequence of errors, which are serially i.i.d. random vectors of dimension $m$ with mean zero, and common joint density $g$. The errors $(\varepsilon_{t})$ are assumed to have a non-Gaussian distribution and are not assumed independent of past $Y$’s. This implies that $(\varepsilon_{t})$ cannot be interpreted as the (causal) innovations.
For the existence of a unique strictly stationary solution to the VAR model (2.1), we assume that the roots of the characteristic equation of the autoregressive polynomial matrix $det(Id-\Phi_{1}\lambda-\cdots\Phi_{p}\lambda^{p})=0$ are of modulus either strictly greater, or strictly smaller than one, i.e. are either outside, or inside the unit circle, and that the joint density $g$ has a uniform tail index $\alpha$. The strictly stationary solution $(Y_{t})$ to model (2.1) can be written as an infinite two-sided moving average (MA($\infty$)) in errors $\varepsilon_{t}$:
@@
This is a linear time series, according to the terminology of Rosenblatt (2012). The autoregressive matrices $\Phi_{1},...,\Phi_{p}$ and the matrices of coefficients $C_{j}$ on the past and future terms of this MA representation are uniquely defined when $\varepsilon_{t}$ is non-Gaussian, which is an identifying assumption. Process $Y_{t}$ is said to be causal in $\varepsilon_{t}$, if $Y_{t}=\sum_{j=0}^{+\infty}C_{j}\varepsilon_{t-j}$, noncausal in $\varepsilon_{t}$, if $Y_{t}=\sum_{j=-\infty}^{-1}C_{j}\varepsilon_{t-j}=\sum_{j=1}^{+\infty}C_{-j}\varepsilon_{t+j}$, or mixed, otherwise.
In the presence of a noncausal component, the assumption of strict stationarity of ($Y_{t}$) implies that, in calendar time, $(Y_{t})$ has a nonlinear dynamic and past-dependent conditional heteroscedasticity. Then, the process $(Y_{t})$ can be characterized by a rather complicated conditional distribution of $Y_{t+h}$ given $\underline{Y_{t}}=(Y_{t},Y_{t-1},...)$ for $h=1,2,..$, that leads us to the nonlinear out-of-sample (oos) predictive distribution described in Section 3. The nonlinearity in the path of $(Y_{t})$ manifests itself through local trends, spikes and bubbles similar to those observed in the time series of commodity (oil) prices, exchange rates, or cryptocurrency prices [Gourieroux and Zakoian (2017), Gourieroux and Jasiak (2017), Gourieroux, Jasiak and Tong (2021)].
@@ 2.2 State-Space Representation
Let us recall the state-space representation of the mixed VAR process with the latent causal and noncausal components of the observed process as state variables.
@@
$\left(\begin{array}[]{c}Y_{t}\\ \tilde{Y}_{t-1}\end{array}\right)\equiv\left(\begin{array}[]{c}Y_{t}\\ Y_{t-1}\\ \vdots\\ Y_{t-p+1}\end{array}\right)=\Psi\left(\begin{array}[]{c}Y_{t-1}\\ \tilde{Y}_{t-2}\end{array}\right)+\left(\begin{array}[]{c}\varepsilon_{t}\\ 0\end{array}\right),\;\;\mbox{with}\;\;\Psi=\left(\begin{array}[]{cccc}\Phi_{1}&\cdots&\cdots&\Phi_{p}\\ Id&0&\cdots&\\ \vdots&\ddots&\ddots&\vdots\\ 0&\cdots&Id&0\end{array}\right).$
The eigenvalues of the autoregressive matrix $\Psi$ are reciprocals of the roots of the characteristic equation (2.4).
@@
Matrix $\Psi$ has a real Jordan representation: $\Psi=A\left(\begin{array}[]{cc}J_{1}&0\\ 0&J_{2}\end{array}\right)A^{-1},$ where $J_{1}$ (resp. $J_{2}$) are real $(n_{1}\times n_{1})$ (resp. $(n_{2}\times n_{2})$) matrices where $n_{2}=n-n_{1}$, with all eigenvalues of modulus strictly less than 1 [resp. strictly larger than 1], and $A$ is a ($n\times n$) invertible matrix [see, Perko (2001), Gourieroux and Jasiak (2017), Section 5.2, for real Jordan representations]. Then, the above equation can be rewritten after the change of basis $A^{-1}$ as:
@@
Let us introduce a block decomposition of $A^{-1}$: $A^{-1}\equiv\left(\begin{array}[]{c}A^{1}\\ A^{2}\end{array}\right),$ where $A^{1}$ is of dimension $(n_{1}\times n)$, and define the transformed variables:
@@
This leads us to two sets of state components of process ($Y_{t}$) such that:
@@
Hence, the state-space representation of process ($Y_{t}$) consists of: state equations (2.5) representing the causal and non-causal dynamics of state variables $Z_{t}$ and measurement equations for $Y_{t}$ obtained by solving:
$\left(\begin{array}[]{c}Y_{t}\\ \tilde{Y}_{t-1}\end{array}\right)=AZ_{t}.$
These measurement equations are deterministic. Therefore, the filtrations generated by processes $(Y_{t})$ and $(Z_{t})$ are identical.
Since the Jordan representation of matrix $\Psi$ is not unique, the state-space representation is not unique either. It depends on the choice of state variables, i.e. the latent factors $Z_{1}$ and $Z_{2}$, up to linear invertible transformations.
c) State-Specific Linear Errors
The first set of state equations in (2.5) defines the causal VAR(1) process ($Z_{1,t}$), which has the causal MA()($\infty$)( ∞ ) representation:
@@
where $\eta_{1,t}$ is a function of $\varepsilon_{t},\varepsilon_{t-1}...$. The second set of state equations in (2.5) needs to be inverted to obtain a MA representation in matrices with eigenvalues of modulus strictly less than 1. We get:
$Z_{2,t}=J_{2}^{-1}Z_{2,t+1}-J_{2}^{-1}\,\eta_{2,t+1}=-\sum_{j=0}^{+\infty}[J_{2}^{-j-1}\,\eta_{2,t+j+1}].$ (2.7)
We observe that $(Z_{2,t})$ is a noncausal process with a one-sided moving average representation in future values $\varepsilon_{t+1},\varepsilon_{t+2},...,$.
Let us now discuss the state-specific linear errors $\eta_{1,t},\eta_{2,t}$. From equations (2.8), (2.9) and the stationarity conditions, it follows that:
@@
where $\underline{Z}_{1,t-1}=(Z_{1,t-1},Z_{1,t-2},...$) and $\bar{Z}_{2,t+1}=(Z_{2,t+1},Z_{2,t+2},...$). Therefore, $\eta_{1,t}$ (resp. $\eta_{2,t}$) can be interpreted as the causal linear innovation of $Z_{1,t}$ based on the information $\underline{Z}_{1,t-1}$ (resp. noncausal linear innovation of $Z_{2,t}$ based on the information $\bar{Z}_{2,t+1}$). It is important to note that the information set $\underline{Z}_{1,t-1}$ (resp. $\bar{Z}_{2,t+1}$) differs in general from the global causal information set $\underline{Y}_{t-1}=\underline{Z}_{t-1}$ (resp $\bar{Y}_{t+1}=\bar{Z}_{t+1}$). In Section 5, we clarify this point by introducing the notion of a nonlinear innovation that takes into account all available information. Before doing that, we need to derive the expressions of forward and backward predictive densities.
3 Out-of-Sample Predictive Density
This Section presents the forecasting and backcasting methods based on closed-form expressions of forward and backward predictive densities.
3.1 Forward Predictive Density
The expression of the predictive density of $Y_{T+1}$ given $\underline{Y}_{T}=(Y_{T},Y_{T-1},...)$ is given below and derived in Online Appendix A.1., assuming that the matrices of autoregressive coefficients $\Phi_{1},\Phi_{2},...$ and joint error density $g$ are known.